# Pluricanonical Geometry of Varieties Isogenous to a Product and Pluricanonical Decompositions of Abelian Covers

### v2

@@ Abstract.
Threefolds of general type, varieties isogenous to a product, finite group actions, abelian covers, canonical map, pluricanonical maps, birationality
@@ Introduction
Next, we give a generalization to higher dimensions and to pluricanonical maps of the representation-theoretic birationality criteria (see Proposition 4.7 and Corollary 4.8) originally introduced by Catanese [Cat18] in the surface case for the canonical map.
@@ 5.1. Example of a Threefold with Birational Canonical Map
Since this quotient map is unramified, the differential $dF$ of the composition $F$ is injective if and only if the $d\Phi_{1}$ is injective. We are going to show that there exists a point where the differential $dF$ is not injective. For this purpose, we work on the affine open set $x_{3}y_{3}z_{3}\neq 0$:
@@
$\begin{split}&y_{0}^{2}=\frac{c^{\prime}(b^{\prime}-c^{\prime})\pm\sqrt{(c^{\prime 2}-4)(b^{\prime}-c^{\prime})^{2}}}{2(b^{\prime}-c^{\prime})^{2}},\quad y_{1}=-\frac{1}{(b^{\prime}-c^{\prime})y_{0}},\quad y_{2}^{2}=-y_{0}^{2}-y_{1}^{2}-a^{\prime}y_{0}y_{1},\quad y_{4}=0\\ &z_{0}^{2}=\frac{c^{\prime\prime}(b^{\prime\prime}-c^{\prime\prime})\pm\sqrt{(c^{\prime\prime 2}-4)(b^{\prime\prime}-c^{\prime\prime})^{2}}}{2(b^{\prime\prime}-c^{\prime\prime})^{2}},\quad z_{1}=-\frac{1}{(b^{\prime\prime}-c^{\prime\prime})z_{0}},\quad z_{2}^{2}=-z_{0}^{2}-z_{1}^{2}-a^{\prime\prime}z_{0}z_{1},\quad z_{4}=0.\end{split}$
@@
(5.2) $J:=\begin{pmatrix}z_{4}&0&x_{0}\\ y_{4}&x_{0}&0\\ \frac{\partial x_{1}}{\partial x_{0}}z_{4}&0&x_{1}\\ \frac{\partial x_{1}}{\partial x_{0}}y_{4}&x_{1}&0\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{0}&x_{2}z_{0}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{0}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{1}&x_{2}z_{1}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{1}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}z_{2}&0&x_{2}\frac{\partial z_{2}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{4}&y_{4}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{0}z_{2}&x_{4}z_{2}\frac{\partial y_{0}}{\partial y_{4}}&x_{4}y_{0}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{1}z_{2}&x_{4}z_{2}\frac{\partial y_{1}}{\partial y_{4}}&x_{4}y_{1}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{2}&x_{4}\frac{\partial y_{2}}{\partial y_{4}}&0\\ \end{pmatrix}$
@@
and thus the first column of the Jacobian matrix of $F$ is zero.
@@
Since $x_{0}\neq 0$ or $x_{1}\neq 0$, this implies that $z_{4}=0.$ We want to show that also $y_{4}=0$. Assume by contradiction $y_{4}\neq 0$. Since either $z_{0}\neq 0$ or $z_{1}\neq 0$, when we consider the minors given by rows 2, $i$ and 12, where $i\in\{8,\dotso,11\}$ and row 12, we obtain that
@@
We are done, since there only finitely many points on
@@ 47 $\mathbb{Z}_{2}^{4}$ 2 2 4 $2^{5}$ $2^{6}$ $2^{6}$ 5 3 0 9 15 0 0 4 0 4 0 4 0
Conversely, let $G\in f^{\ast}\operatorname{Nef}(Y)$ be a $\mathbb{Q}$-divisor. By projection formula, $G\cdot C=0$ for each curve $C$ contracted by $f$. Suppose by contradiction that $G\notin\mathcal{F}$. Since $f\colon X\to Y$ is not an isomorphism, it contracts some curves, hence $G\notin\operatorname{Amp}(X)$. This implies that there exists an extremal face $\mathcal{F}^{\prime}\neq\mathcal{F}$ such that $G\in\mathcal{F}^{\prime}$. Thus, there exists $H\in\operatorname{Amp}(X)$ such that $H\in(D,G)$. By the same argument given in the proof of Claim 2 we conclude that $H\cdot C=0$ for each curve $C$ contracted by $f$, a contradiction. Therefore, $G\in\mathcal{F}$. ∎

### v3

# arXiv:2512.21294v3[math.AG] 02 Mar 2026

@@ Abstract.
Abelian covers, finite group actions, canonical map, pluricanonical maps, birationality, varieties isogenous to a product, threefolds of general type
@@ Introduction
Next, we give a generalization to higher dimensions and to pluricanonical maps of the representation-theoretic birationality criteria (see Proposition 4.5, Corollary 4.6, Proposition 4.7 and Corollary 4.8) originally introduced by Catanese [Cat18] in the surface case for the canonical map.
@@ 5.1. Example of a Threefold with Birational Canonical Map
Since this quotient map is unramified, the differential $dF$ of the composition $F$ is injective if and only if $d\Phi_{1}$ is injective. We are going to show that there exists a point where the differential $dF$ is not injective. For this purpose, we work on the affine open set $x_{3}y_{3}z_{3}\neq 0$:
@@
$\begin{split}&y_{0}^{2}=\frac{c^{\prime}(b^{\prime}-c^{\prime})\pm\sqrt{(c^{\prime 2}-4)(b^{\prime}-c^{\prime})^{2}}}{2(b^{\prime}-c^{\prime})^{2}},\quad y_{1}=-\frac{1}{(b^{\prime}-c^{\prime})y_{0}},\quad y_{2}^{2}=-y_{0}^{2}-y_{1}^{2}-a^{\prime}y_{0}y_{1},\quad y_{4}=0,\\ &z_{0}^{2}=\frac{c^{\prime\prime}(b^{\prime\prime}-c^{\prime\prime})\pm\sqrt{(c^{\prime\prime 2}-4)(b^{\prime\prime}-c^{\prime\prime})^{2}}}{2(b^{\prime\prime}-c^{\prime\prime})^{2}},\quad z_{1}=-\frac{1}{(b^{\prime\prime}-c^{\prime\prime})z_{0}},\quad z_{2}^{2}=-z_{0}^{2}-z_{1}^{2}-a^{\prime\prime}z_{0}z_{1},\quad z_{4}=0.\end{split}$
@@
$J:=\begin{pmatrix}z_{4}&0&x_{0}\\ y_{4}&x_{0}&0\\ \frac{\partial x_{1}}{\partial x_{0}}z_{4}&0&x_{1}\\ \frac{\partial x_{1}}{\partial x_{0}}y_{4}&x_{1}&0\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{0}&x_{2}z_{0}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{0}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}y_{2}z_{1}&x_{2}z_{1}\frac{\partial y_{2}}{\partial y_{4}}&x_{2}y_{2}\frac{\partial z_{1}}{\partial z_{4}}\\ \frac{\partial x_{2}}{\partial x_{0}}z_{2}&0&x_{2}\frac{\partial z_{2}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{0}}{\partial y_{4}}&y_{0}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{0}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{0}}{\partial z_{4}}\\ 0&z_{1}\frac{\partial y_{1}}{\partial y_{4}}&y_{1}\frac{\partial z_{1}}{\partial z_{4}}\\ 0&z_{4}&y_{4}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{0}z_{2}&x_{4}z_{2}\frac{\partial y_{0}}{\partial y_{4}}&x_{4}y_{0}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{1}z_{2}&x_{4}z_{2}\frac{\partial y_{1}}{\partial y_{4}}&x_{4}y_{1}\frac{\partial z_{2}}{\partial z_{4}}\\ \frac{\partial x_{4}}{\partial x_{0}}y_{2}&x_{4}\frac{\partial y_{2}}{\partial y_{4}}&0\\ \end{pmatrix}$
@@
and thus the first column of the Jacobian matrix $J$ is zero.
@@
Since $x_{0}\neq 0$ or $x_{1}\neq 0$, this implies $z_{4}=0.$ We want to show that also $y_{4}=0$. Assume by contradiction $y_{4}\neq 0$. Since either $z_{0}\neq 0$ or $z_{1}\neq 0$, when we consider all the minors given by row 2, row $i$ with $i\in\{8,\dotso,11\}$ and row 12, we obtain that
@@
We are done, since there are only finitely many points on
@@ 47 $\mathbb{Z}_{2}^{4}$ 2 2 4 $2^{5}$ $2^{6}$ $2^{6}$ 5 3 0 9 15 0 0 4 0 4 0 4 0
Conversely, let $G\in f^{\ast}\operatorname{Nef}(Y)$ be a $\mathbb{Q}$-divisor. By projection formula, $G\cdot C=0$ for each curve $C$ contracted by $f$, i.e., $D\cdot C=0$. Suppose by contradiction $G\notin\mathcal{F}$. Since $f\colon X\to Y$ is not an isomorphism, it contracts some curves, hence $G\notin\operatorname{Amp}(X)$. This implies that there exists an extremal face $\mathcal{F}^{\prime}\neq\mathcal{F}$ such that $G\in\mathcal{F}^{\prime}$. Suppose first $\dim\mathcal{F}=2$ and assume w.l.o.g. $\mathcal{F}=\mathbb{R}_{\geq 0}F_{1}+\mathbb{R}_{\geq 0}F_{2}$. Hence, it must occur
$\mathcal{F}^{\prime}\in\{\mathbb{R}_{\geq 0}F_{3},\quad\mathbb{R}_{\geq 0}F_{2}+\mathbb{R}_{\geq 0}F_{3},\quad\mathbb{R}_{\geq 0}F_{1}+\mathbb{R}_{\geq 0}F_{3}\}.$
Let $C^{\prime}:=F_{1}\cdot F_{2}$. One can easily verify that in each of these cases $D\cdot C^{\prime}=0$, but $G\cdot C^{\prime}>0$, a contradiction. Suppose finally $\dim\mathcal{F}=1$ and assume w.l.o.g. $\mathcal{F}=\mathbb{R}_{\geq 0}F_{1}$. Hence, it must occur
$\mathcal{F}^{\prime}\in\{\mathbb{R}_{\geq 0}F_{2},\quad\mathbb{R}_{\geq 0}F_{1}+\mathbb{R}_{\geq 0}F_{2},\quad\mathbb{R}_{\geq 0}F_{2}+\mathbb{R}_{\geq 0}F_{3},\quad\mathbb{R}_{\geq 0}F_{3},\quad\mathbb{R}_{\geq 0}F_{1}+\mathbb{R}_{\geq 0}F_{3}\}.$
Taking respectively $C^{\prime}:=F_{1}\cdot F_{3}$ in the first three cases and $C^{\prime}:=F_{1}\cdot F_{2}$ in the last two, we get as before $D\cdot C^{\prime}=0$, but $G\cdot C^{\prime}>0$, again a contradiction. Therefore, $G\in\mathcal{F}$. ∎